TaskMaster Enterprises employs a standard cost system in which direct materials inventory is carried at standard cost. TaskMaster has established the following standards for the prime costs of one unit of product.

Standard Standard Standard
Quantity Price Cost
Direct Materials 8 pounds $ 1.80 per pound $ 14.40
Direct Labor 0.25 hour $ 8.00 per hour 2.00
$ 16.40

During November, TaskMaster purchased 160,000 pounds of direct materials at a total cost of $304,000. The total factory wages for November were $42,000, 90% of which were for direct labor. TaskMaster manufactured 19,000 units of product during November using 142,500 pounds of direct materials and 5,000 direct labor hours.

Is the direct materials efficiency variance favorable or unfavorable?

Multiple Choice

  • Favorable

  • Unfavorable
Answer #1

Materials efficiency variance

= (SQ-AQ) *SP

= (19,000*8 - 142,500)*1.80

= (152,000 - 142,500)*1.80

FAVOURABLE